(1 point) Find the solution to initial value problem ,dy – 14 + 49y = 0, y(0) = 2, y(0) = 3 dt g(t) =
Answer #1

d dt o lu doll + ugy zo , lo) 22, yCO)=3 Ausilary equatis is r²_14r +49=0 e²-74-7 2+49=0 =) rCr-7)=7 (5-7)=0 > (5-7)²=0 r